Status Corporation purchased for $350,000 a new numerical controller during the last month of 2012. Extra installation costs were $40,000. The recovery period was 7 years with an estimated salvage value of 10% of the original purchase price. Status sold the system at the end of 2015 for $45,000. State the numerical values for the following: remaining life at sale time, market value in 2015, book value at sale time if 65% of the unadjusted basis had been depreciated.

Question 13 options:

Remaining life = 7 years

Market value = $305,000

Book value = $35,000

Remaining life = 3 years

Market value = $310,000

Book value = $45,000

Remaining life = 4 years

Market value = $45,000

Book value = $136,500

Remaining life = 4 years

Market value = $40,000

Book value = $253,500
Answer #1

Original Useful life of numerical controller = 7 years

It was purchased in the last month of 2012 and sold at the end of 2015. Thus, it was used for 3 years.

Remaining life of numerical controller = 7 - 3

= 4 years

Status sold the system at the end of 2015 for $45,000.

Thus, market value = $45,000

Book value = Cost price - Accumulated depreciation

= 390,000 - 390,000 x 65%

= 390,000 - 253,500

= $136,500

Correct option is (c)

Kindly comment if you need further assistance. Thanks